John is twice as old as Sally and Sally is three times as old as Mary. There total age is 100 years. How old is each person?
Answer by TimothyLamb(4379)

x = john age
y = sally age
z = mary age
---
x = 2y
y = 3z
x + y + z = 100
---
put the system of linear equations into standard form
---
x - 2y = 0
y - 3z = 0
x + y + z = 100
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = john age = 60
y = sally age = 30
z = mary age = 10
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php